OpenStudy (anonymous):

anti derivative of ((x - 1)^2) / x^2

OpenStudy (jamesj):

Note first that ((x - 1)^2) / x^2 = ( x^2 - 2x + 1 ) / x^2 = 1 - 2/x + 1/x^2 Can you integrate it now?

OpenStudy (anonymous):

=( x^2 - 2x + 1 )/ x^2 = 1 - 2/x + 1/x^2 integrate these one by one

OpenStudy (anonymous):

yes thank you, I did the first step wrong!